OpenStudy (anonymous):

Write an equation in slope intercept form of the line that passes through (4, 1) and (5, -1). Answer y = -4x + 1 y = -1/2 x + 9 y = 4x + 1 y = -2x + 9 1 points Question 8 Write an equation of the line that passes through (8, 3) and (4, 3). Answer y = 3 y = 4x + 3 y = 4 y = 3x

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(4\quad ,&1)\quad &(5\quad ,&-1) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1} \\ \quad \\ y-y_1=m(x-x_1)\qquad \textit{plug in your values and solve for "y"}\) \(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(8\quad ,&3)\quad &(4\quad ,&3) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1} \\ \quad \\ y-y_1=m(x-x_1)\qquad \textit{plug in your values and solve for "y"}\)